OpenStudy (anonymous):

etermine the three cubic roots of 8.give the answer in the form of a+bj.also give the main root in the exponential form

OpenStudy (anonymous):

a^3-b^3 = (a-b)(a^2+ab+b^2)

OpenStudy (anonymous):

Can you please do this sum I don't understand clearly

OpenStudy (anonymous):

x^3=8

OpenStudy (anonymous):

where a^3=x^3 and b^3 = 8 solve for a and b and plug into the first equation given,

OpenStudy (anonymous):

I must write 8 in exponential form first or ?

OpenStudy (anonymous):

well it wouldnt hurt

OpenStudy (anonymous):

Ok,just do it in your way so that I can compare it with myn wanna see whre I made mistakes please its an assignment

OpenStudy (anonymous):

a=? b=?

OpenStudy (anonymous):

a=2 b=3

OpenStudy (anonymous):

ok so x^3=8 can also be written as x^3-8=0 and can be factored by using the equations i provided by substituting in the values you solved for

OpenStudy (anonymous):

x^3=a^3 x=a 8=b^3 b=2